OpenStudy (goho2838):

Use the distributive property to simplify the expression 4 × 24.

Directrix (directrix):

Using the Distributive Property (see attachment), I assume you are to do this: 4 x 24 = 4*(20 + 4) = 4*20 + 4*4 = ? + ? = ? @goho2838

Directrix (directrix):

@goho2838 Please post what goes in the blanks so that we can be sure you cranked out the problem correctly, okay?